I Need to put (o,3) and (2,-3) in an equation in standard form. Please help!
Answer by stanbon(75887)

I Need to put (o,3) and (2,-3) in an equation in standard form.
----
slope =(3--3)/(0-2) = 6/-2 = -3
---------------
Form:: y = mx + b
-----
b = f(0) = 3
-------------
Ans: y = -3x + 3
-------------------
Cheers,
Stan H.
--------------